OpenStudy (anonymous):

How to do epsilon-delta proof of lim_(x->0) (x^2-2x+3)=3?

OpenStudy (anonymous):

http://www.classicshorts.com/stories/cwrdmaup.html

OpenStudy (anonymous):

?

OpenStudy (anonymous):

sorry. that last response was supposed to go to somebody else, but for some reason got put here. sorry about that!

OpenStudy (anonymous):

uaheuhuheuaehhhaeuhuehuhe

OpenStudy (anonymous):

i gave you answer once already

OpenStudy (experimentx):

at x's tends to 0, all terms of function tends to 0 .... except constant 3, so 3

OpenStudy (anonymous):

No, you did not.

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

I know that: 0<|x - 0| < delta, so |f(x) -3|< epsilon

OpenStudy (anonymous):

|x^2 -2x +3 - 3| < epsilon

OpenStudy (anonymous):

|x^2 - 2x| < epsilon |x(x-2)| = |x| |x-2| = epsilon C Taking |x-2| < C epsilon/C = delta And now what i do?

OpenStudy (experimentx):

|x| |x-2| < e => |x - 0| < e / |x-2| < d (suppose it) => 3 is limit as x->0

OpenStudy (anonymous):

Sorry, I did not understand.

OpenStudy (experimentx):

beginning with |x^2 -2x +3 - 3| < epsilon show |x - 0| < d

OpenStudy (anonymous):

I need a answer with more formalization.